OpenStudy (anonymous):

If 3.8 moles of zinc metal react with 6.5 moles of silver nitrate, how many moles of silver metal can be formed, and how many moles of the excess reactant will be left over when the reaction is complete?

OpenStudy (anonymous):

Unbalanced equation: Zn + AgNO3 → Zn(NO3)2 + Ag

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@jadaaaa, at first you should balanced that equation.

OpenStudy (anonymous):

So their molar ratio is 1:2 - 1:2 Now according with your values, if you have 3.8 mol of Zn, this should react with 3.8*2=7.6 mol of AgNO3, not 6.5, telling you that Zn is in excess. Hence if we have only 6.5mol of AgNO3, it would completely react with 6.5/2=3.25mol of Zn, meaning 3.8-3.25=0.55 mol are in excess. As you can see in the molar ration of reactants and products, AgNO3 and Ag have the same molar ratio of 2:2 or 1:1, therefore6.5 mol of Ag are formed.

OpenStudy (anonymous):

Thank you, I understand it a little better now. @Hoslos